# FilterPrompt: Guiding Image Transfer in Diffusion Models

### v2

@@ Abstract
In controllable generation tasks, flexibly manipulating the generated images to attain a desired appearance or structure based on a single input image cue remains a critical and longstanding challenge. Achieving this requires the effective decoupling of key attributes within the input image data, aiming to get representations accurately. Previous research has predominantly concentrated on disentangling image attributes within feature space. However, the complex distribution present in real-world data often makes the application of such decoupling algorithms to other datasets challenging. Moreover, the granularity of control over feature encoding frequently fails to meet specific task requirements. Upon scrutinizing the characteristics of various generative models, we have observed that the input sensitivity and dynamic evolution properties of the diffusion model can be effectively fused with the explicit decomposition operation in pixel space. This integration enables the image processing operations performed in pixel space for a specific feature distribution of the input image, and can achieve the desired control effect in the generated results. Therefore, we propose FilterPrompt, an approach to enhance the model control effect. It can be universally applied to any diffusion model, allowing users to adjust the representation of specific image features in accordance with task requirements, thereby facilitating more precise and controllable generation outcomes. In particular, our designed experiments demonstrate that the FilterPrompt optimizes feature correlation, mitigates content conflicts during the generation process, and enhances the model’s control capability.
Keywords:
Visual Prompt, Image Transfer, Image-to-Image Translation, Explicit Decomposition, Controllable Generation, Diffusion Models
Figure 1: Comparison of image transfer tasks. Our approach FilterPrompt enables image transfer in multiple domains at local, object-centric, and full-graph levels. Compared to other works [1, 50, 52], our approach can help the model better preserve the geometric properties of structural images while maintaining consistent color distribution and texture features with appearance images.
@@ 1 Introduction
In the classic realm of Controllable Image Generation, achieving flexible control over the characteristics, attributes, or style of generated images remains a focal point of research. Researchers typically focus on acquiring low-dimensional feature representations of high-dimensional image features while minimizing the issues of information loss and attribute entanglement [5]. Alternatively, they seek solutions that offer greater interpretability in the generation process, promoting a more stable and efficient model training process, thus fostering further advancements in the field of image generation.
@@
Controllable generation typically follows two approaches: 1. Some works involve disentangling the characteristics of an input image in the feature space and obtaining feature representations relevant to the control objective. Subsequently, the network regulates the degree to which these feature representations are expressed in the generated image, employing a diverse array of meticulously designed loss metrics [8, 4, 51, 29, 46, 16]. 2. Another category of work involves incorporating a conditioning mechanism into the architecture of the model. These works improve the capacity of the model to integrate control conditions while learning the target style domain’s data distribution. Consequently, the generated image exhibits a style map akin to the training set, yielding an artistic effect that aligns with the input content image [34, 21, 30, 43].
@@
The above controllable generation ideas have their limitations. On the one hand, there will be an unavoidable loss when mapping different data domains to the same feature space. Simultaneously, it is impossible to ensure that there is no entanglement of attributes among the representations [48, 6]. On the other hand, using machine learning algorithms to train style mappings may not have the same level of interpretability as traditional mathematical modeling approaches, and the training of such models often requires expensive datasets [7, 45].
Hence, unlike previous endeavors focused on refining algorithms and models within the feature space, we redirect our focus toward the pixel space. Intuitively, certain semantic features, often indistinguishable in models, exhibit discernible distribution discrepancies visible to the naked eye in the pixel space. Moreover, manipulating abstract feature representations within the feature space may inadvertently impact the entire result while executing image processing operations on input data within the pixel space is more intuitive and controllable.
Figure 2: Enhance network performance with filtering operations. Filtering operations can enhance image quality. When neural networks extract specific image features, strategically incorporating filtering operations upstream enables targeted suppression or enhancement of particular feature distributions.
@@
Following experimental comparisons of various mainstream generative models, we observe that the diffusion model possesses properties of input sensitivity and dynamic evolution which are very sort for image processing operations in the pixel space. This integration enables image processing operations performed in pixel space for a specific feature distribution of the input image and can achieve the desired control effect in the generated results. Therefore, we name this approach FilterPrompt.
@@
To validate the aforementioned assertion, we build a task framework for experimental testing based on the existing pre-trained model and use it to demonstrate the impact of FilterPrompt on the control effect of appearance features and structure features of the generated image. Finally, we conduct quantitative analyses of the generated results. These analyses demonstrate that the FilterPrompt optimizes feature correlation, mitigates content conflicts during the generation process, and enhances the model’s control capability. We believe that our approach can be easily generalized to more complex combinations. At the same time, FilterPrompt has the characteristics of requiring no additional training and being highly interpretable.
@@
In summary, our contributions are shown as follows: 1. We introduce a new approach, FilterPrompt, aimed at optimizing model control effects. Our approach can be customized for various input data based on specific task requirements and combined with the diffusion model to achieve the expected control effect. 2. We analyze how FilterPrompt facilitates the desired control effect within the diffusion model framework. Additionally, we have designed experiments to provide explanations to demonstrate the feasibility of our approach. 3. Our experiments encompass a range of tasks, including local, object-centered, and full-graph appearance transfer, as well as image style transfer. We present the application of FilterPrompt in these tasks and compare its performance with various other models. Through experimentation, we substantiate the efficacy of our approach in image transfer.
@@ 2 Related Work
2.1 Controllable Image Generation in Diffusion Modeling
In the development of the diffusion model, the research on the Controllable Image Generation task has roughly three stages. The first stage of the work is mainly based on the iterative denoising process and achieves controllable generation effects by rationally using the input image to generate a deterministic guided generation paradigm [9, 33, 31]. This stage of the work controls the semantic similarity between the generated image and the input image by influencing the proportion of information mixing in the denoising network of U-Net. The second phase of the work is based on the image generation strategy guided by the display classifier. Optimization is performed by adding gradient information from the classifier to the loss function. This idea first originated from the concept of "Classifier Guidance," proposed by[42] and further advanced in[11, 37]. Since then, numerous studies have broadened the scope of classifiers, extending the classification guidance of the diffusion model to encompass diverse modalities such as text, images, and other multi-modal data [2, 22, 47, 16, 28]. The third stage marks the era of large models based on implicit classifiers. To address the issue of declining diversity in Classifier Guidance, the "Classifier-Free Guidance" strategy [18] emerges later. This approach involves decomposing the gradient guidance from the explicit classifier into two components. One component is an unconditionally generated gradient prediction model, akin to the conventional DDPM [17]. The other is a gradient prediction model based on conditional generation, conceptualized as a U-Net network with an overlay of a cross-attention mechanism. The success of this approach has catalyzed the evolution of a variety of subsequent image editing technologies. These technologies utilize the diffusion model as a foundational framework and integrate the attention mechanism, resulting in notable progress in the application of the diffusion model across diverse fields. Notable projects in this domain include DALLE-2 [36], DreamFusion [35], Stable Diffusion Model (SDM) [38], and more.
Here, our primary focus lies on Grounded Generation and Layout-driven Generation within the context of the controllable generation problem. Representative works in this area include GLIGEN [26], ControlNet [52], and IP-Adapter [50]. We aim to investigate the generative capabilities of diffusion models in addressing semantic-level conditional guidance, particularly in scenarios with limited sample sizes.
@@ 2.2 Explicit Decomposition
Traditionally, filtering algorithms have been considered the explicit decomposition approach as they can break down input data into components at different frequencies or spatial scales. Examples of common filters include the Gaussian filter, Sobel filter [41], Adaptive filter [20], and Gabor filter [12, 39]. These filters can weigh data at different frequencies or scales to suppress or enhance specific features, as illustrated in Figure 2. Performing preprocessing operations helps neural networks better understand the structure and features of data in order to acquire more refined representations.
@@ 3 Prerequisite
3.1 Prompt Impact on the Iterative Process of Diffusion Models
Specifically, the forward diffusion defines a known Gaussian translation process. So the image intermediate quantity $x_{t}$ at each moment of the forward process can establish a unique relationship with the input image $x_{0}$ as:
@@ 3.1 Prompt Impact on the Iterative Process of Diffusion Models
The focus of this process is the network’s prediction of noise distribution: $\epsilon_{\theta}\left(x_{t},c,t\right)$. This prediction is affected by the current moment $x_{t}$ and condition $c$, and condition $c$ comes from the external reference image $y$ mapping result. Therefore, if we perform filters $f_{\gamma}$ on either side of the input like making $c=\mathrm{Encoder}(y)$ become ${c{{}^{\prime}}=~{}\mathrm{Encoder}(f_{\gamma}(y))~{}}$, it will affect the prediction results of the noise distribution $\epsilon_{\theta}$, and even affect the migration direction of the generated distribution at that time node as illustrated in Figure 3. The impact caused by conditions $c$ or the input content image $x_{0}$ would be reiterated at each sampling instance, and the minor changes introduced by filters would be involved in the entire generation process.
In the subsequent experiment depicted in Figure 8, we observe that, after applying filtering operations to a specific feature distribution in pixel space, the degree of expression for that feature distribution aligns with expectations. Moreover, the filtering operations targeting a specific feature distribution does not affect the expression of other features. This indicates that the influence of filters is independent in the feature space from the encoding of other representations. It can be seen that prompts in pixel space offer a lightweight and convenient way to optimize the entanglement between feature representations in the diffusion model.
@@
Figure 3: Part (a) shows the illustration of filter’s impact on the corresponding sampling inference stages in the static generative model and the dynamic generative model. Part (b) gives a comparison of the results obtained by applying filter to some works of generative models. These include FastStylePredict [15], which is a fast style transfer algorithm based on CNN, CycleGAN [55], DualStyleGAN [49] based on GAN and AE architecture, InST [53] and IP-Adapter [50] based on SDM architecture and the effects of our work. Comparing the last column of figures, we can intuitively see that filter operation has a more significant impact on diffusion models.
@@
In the early days, the fast style transfer algorithm [14, 25, 15] is based on convolutional neural network (CNN), AutoEncoder (AE), and Generative Adversarial Network (GAN) [10, 55, 49], and their fitting effect on the target data distribution depends on the design of the training process. During the training process, the model attempts to capture the statistical characteristics of the entire data distribution and obtain the maximum likelihood representation of the data distribution. When training is completed, the generated samples exist statically in the latent space. Therefore, we consider this type of model to be the static generative model.
@@
We noticed that prompts in pixel space may exhibit certain limitations in previous static generation models, but it can be well combined with dynamic generation models. In the style transfer task shown in Figure 3, We applied the same level of blur interference to the input data of various generative models. The results indicate that the impact of blur prompt on static generative models is relatively weak, while the outcomes of diffusion models show significant changes. We believe that for static generation models, using prompt on input data during the sampling inference stage will only have a slight impact on the mapping position of this input in the static distribution. And this position is close to the mapping result of the original image $x$. Therefore, prompts in pixel space will not significantly affect the generation results of static generation models. However, for diffusion models, as indicated by Equation 1, there is a predefined long-term dependency between the current time node image $x_{t}$ and the input image $x_{0}$, in the forward process. Therefore, the impact of using prompts on the input content image $x_{0}$ will be executed again at each sampling, and the minor changes caused by prompts will participate in the entire generation process, thereby exerting a more significant influence on the generation results.
4 Proposed Approach
@@
It is noted that distinguishing and processing analogous feature representations in feature space leads to significant computational overhead. Consequently, we propose the adoption of our pixel-space methodology, termed FilterPrompt. This approach directly manipulates the frequency or distribution characteristics of specific image attributes, thereby influencing the subsequent expression levels of the representation. FilterPrompt offers an intuitive and straightforward approach. Fundamentally, within the diffusion model framework, our approach facilitates deterministic guidance on the generation process through the integration of filtering operations.
To provide a more detailed exposition and validation of the aforementioned statement, the arrangement of this chapter is as follows: In Section 4.1 We provided a general description of the FilterPrompt. In Section 4.2 To verify and discuss the impact of our approach on the control effect of diffusion models, we provide the experimental setup, including the construction of data sets and the framework of baseline. In Section 4.3 we give the effects of various FilterPrompts on the baseline. In Section 4.4 we show the visual explanation of FilterPrompt on the results of each sampling in the denoising generation process.

### v3

# FilterPrompt: A Simple yet Efficient Approach to Guide Image Appearance Transfer in Diffusion Models

@@ Abstract
In controllable generation tasks, flexibly manipulating the generated images to attain a desired appearance or structure based on a single input image cue remains a critical and longstanding challenge. Achieving this requires the effective decoupling of key attributes within the input image data to achieve representations accurately. Previous works have concentrated predominantly on disentangling image attributes within feature space. However, the complex distribution present in real-world data often makes the application of such decoupling algorithms to other datasets challenging. Moreover, the granularity of control over feature encoding frequently fails to meet specific task requirements. Upon scrutinizing the characteristics of various generative models, we have observed that the input sensitivity and dynamic evolution properties of the diffusion model can be effectively fused with the explicit decomposition operation in pixel space. This allows the operation that we design and use in pixel space to achieve the desired control effect on the specific representation in the generated results. Therefore, we propose FilterPrompt, an approach to enhance the effect of controllable generation. It can be universally applied to any diffusion model, allowing users to adjust the representation of specific image features in accordance with task requirements, thereby facilitating more precise and controllable generation outcomes. In particular, our designed experiments demonstrate that the FilterPrompt optimizes feature correlation, mitigates content conflicts during the generation process, and enhances the effect of controllable generation, as shown in Figure 1.
Figure 1: The comparison of generated results. Our approach FilterPrompt enables appearance transfer in multiple domains at local, object-centric, and full-graph levels. Compared to previous works like Cross-Image [1], IP-Adapter [56] and baseline (IP-adapter [56] +ControlNet [58]), our approach can help the model better preserve the geometric properties of structural images while maintaining consistent color distribution and texture features with appearance images.
@@ 1 Introduction
In controllable image generation, achieving flexible control over the appearance attributes of objects in the generated images, such as texture and material, remains a research focus [5]. Some researchers concentrate on refining the data, aiming to acquire the low-dimensional feature representations of input images [39, 31, 53]. Concurrently, another faction of researchers is interested in improving the model architecture. They employ deep learning techniques such as autoencoders (AE), variational autoencoders (VAE), and generative adversarial networks (GAN) to fine-tune feature extraction and processing methods, thereby enhancing the capacity of models to handle complex data autonomously [18, 27, 2, 14].
@@
Specifically, controllable generation typically follows two ways: First, disentangling the characteristics of an input image in the feature space and obtaining feature representations relevant to the control objective. Subsequently, the network regulates the degree to which these feature representations are expressed in the generated image, employing a diverse array of meticulously designed loss metrics [8, 4, 57, 31, 51, 17]. Second, involving incorporating a conditioning mechanism into the architecture of the model. These works improve the capacity of the model to integrate control conditions while learning the target domain’s data distribution [37, 23, 33, 48]. Then, the generated images exhibit an artistic effect similar to the appearance of the training data set and consistent with the input image structure.
@@
However, the above controllable generation ideas have their limitations. On the one hand, mapping different data domains to the same feature space will incur high computational costs. Information loss in this process cannot be avoided while there is also the problem of attribute entanglement between the representations obtained in the feature space [54, 6]. On the other hand, using machine learning algorithms to train style mappings may not have the same level of interpretability as traditional mathematical modeling approaches, and the training of such models often requires expensive data collection [7, 50].
@@
Hence, unlike previous endeavors that focused on refining algorithms and models within the feature space, we redirect our focus toward the pixel space. Intuitively, certain semantic features, often indistinguishable in models, exhibit discernible distribution discrepancies visible to the naked eye in the pixel space. Following experimental comparisons of various mainstream generative models, we observe that the diffusion model possesses properties of input sensitivity and dynamic evolution which are very suitable for image processing operations in the pixel space. This integration enables image processing operations performed in the pixel space for a specific feature distribution of the input image and can achieve the desired control effect in the generated results. Therefore, we name this approach FilterPrompt as shown in Figure 2.
Figure 2: Our FilterPrompt. When diffusion models extract image features, strategically incorporating filtering operations enables targeted suppression or enhancement of particular feature distributions. The filters enhance the performance of diffusion models to improve the quality of generated images.
@@
To validate the aforementioned statement, we build a framework for experimental testing based on the existing pre-trained model and use it to demonstrate the impact of FilterPrompt on the control effect of appearance features and structure features of the generated image. Then, we conduct quantitative analyses of the generated results. These analyses demonstrate that the FilterPrompt optimizes feature correlation, mitigates content conflicts during the generation process, and enhances the model’s control capability. At the same time, our approach can be easily generalized to more complex combinations without requiring additional training and is highly interpretable.
@@
In summary, our contributions are as follows: 1. We introduce a new approach, FilterPrompt, aimed at optimizing model control effects. Our approach can be customized for various input data based on specific task requirements and combined with the diffusion model to achieve the expected control effect. 2. We analyze how FilterPrompt facilitates the desired control effect within the diffusion model framework. Additionally, we have designed experiments to provide explanations to demonstrate the feasibility of our approach. 3. Our experiments encompass a range of tasks for the local, object-centered, and full-graph appearance transfer. We present the application of FilterPrompt in these tasks and compare its performance with various other models. Through experimentation, we substantiate the efficacy of our approach in image transfer.
@@ 2 Related Work
2.1 Controllable Generation in Diffusion Modeling
The research on the controllable image generation task in the diffusion model can be broadly divided into three stages.
The first stage is mainly based on the iterative denoising process and achieves controllable generation effects by rationally using the input image to generate a deterministic guided generation paradigm [9, 36, 34]. This stage of the work controls the semantic similarity between the generated image and the input image by influencing the proportion of information mixing in the denoising network of U-Net. The second stage is based on the image generation strategy guided by the display classifier. Optimization is performed by adding gradient information from the classifier to the loss function. This idea first originated from classifier guidance[46] and further advanced in[11, 42]. Since then, numerous studies have broadened the scope of classifiers, extending the classification guidance of the diffusion model to encompass diverse modalities such as text, images, and other multi-modal data [3, 24, 52, 17, 32]. The third stage marks the era of large models based on implicit classifiers. To address the issue of declining diversity in classifier guidance, classifier-free guidance strategy [19] emerges later. This approach involves decomposing the gradient guidance from the explicit classifier into two components. One component is an unconditionally generated gradient prediction model, akin to the conventional DDPM [20]. The other is a gradient prediction model based on conditional generation, conceptualized as a U-Net network with an overlay of a cross-attention mechanism. The success of this approach has catalyzed the evolution of a variety of subsequent image editing technologies. These technologies utilize the diffusion model as a foundational framework and integrate the attention mechanism, resulting in notable progress in the application of the diffusion model across diverse fields. Notable projects in this domain include DALLE-2 [40], DreamFusion [38], Stable Diffusion Model (SDM) [41], and more.
Here, our primary focus lies on Grounded Generation and Layout-driven Generation within the context of the controllable generation problem. Representative works in this area include GLIGEN [29], ControlNet [58], and IP-Adapter [56]. We aim to investigate the generative capabilities of diffusion models in addressing semantic-level conditional guidance, particularly in scenarios with limited sample sizes. For example, appearance transfer task [47]. It needs to preserve the structure of the target image while applying the desired appearance attributes.
@@ 2.2 Explicit Decomposition
Traditionally, filtering algorithms have been considered the explicit decomposition approach as they can break down input data into components at different frequencies or spatial scales. Examples of common filters include the Gaussian filter, Sobel filter [45], Adaptive filter [22], and Gabor filter [12, 43]. These filters can weigh data at different frequencies or scales to suppress or enhance specific features. Therefore, performing preprocessing operations may help neural networks better understand the structure and features of data in order to acquire more refined representations.
@@ 3 Prerequisite
Figure 3: Filter impact on sampling inference process. After applying a Gaussian filter, the underlying texture in the sampled images changes from a distribution resembling arc patterns to a point-like distribution. Additionally, as shown in the enlarged illustration on the right, it is evident that the use of filters consistently disrupts the expression of redundant pattern features.
3.1 Prompt Impact on Diffusion Models
The forward diffusion defines a known Gaussian translation process. Then the image intermediate quantity $x_{t}$ at each moment of the forward process can establish a unique relationship with the input image $x_{0}$ as:
@@ 3.1 Prompt Impact on Diffusion Models
The focus of this process is the network’s prediction of noise distribution: $\epsilon_{\theta}\left(x_{t},c,t\right)$. This prediction is affected by the current moment $x_{t}$ and condition $c$, and condition $c$ comes from the external reference image $y$ mapping result. Therefore, if we perform filters $f_{\gamma}$ on either side of the input like making $c=\mathrm{Encoder}(y)$ become ${c{{}^{\prime}}=~{}\mathrm{Encoder}(f_{\gamma}(y))~{}}$, it will affect the prediction results of the noise distribution $\epsilon_{\theta}$, and even affect the migration direction of the generated distribution at that time node as illustrated in Figure 4. The impact caused by conditions $c$ or the input structure image $x_{0}$ would be reiterated at each sampling instance, and the minor changes introduced by filters would be involved in the entire generation process.
We observe that after applying filtering operations to a specific feature distribution in pixel space, the degree of expression for that distribution aligns with expectations. Moreover, the filtering operations targeting a specific feature distribution do not affect the expression of other features. This indicates that the influence of filters is independent in the feature space from the encoding of other representations. Therefore, prompts in pixel space offer a lightweight and convenient way to optimize the entanglement between feature representations in the diffusion model.
Based on the framework of combining ControlNet [58] and IP-Adapter [56], we examine the impact of filters at various stages of the sampling inference process and elucidate that the filter plays a guiding role in guiding the Gaussian distribution of the current data toward the target distribution during the migration and diffusion process.
We perform a detailed analysis of the sampling inference stage for the task of converting a bronze sketch to a photo, as shown in Figure 3. The sampled results vividly illustrate that detailed representations of arc patterns initially present in the early stages are weakened by the Gaussian filter, manifesting as point-like distributions. Besides, the Gaussian filter disrupts the continuous semantic expression of patterns. Concurrently, subsequent texture generation doesn’t emerge the negative impact of full-graph blurring, so it illustrates the property that the filtering operation is only effective for specific feature distributions.
The third row of the evolution sequence in Figure 3 illustrates the impact of applying the Gaussian filter only in the first 12 steps, where we observe the absence of redundant pattern features in the final generation result. A set of comparisons in lines 4 and 5 also showcase the effectiveness. In the first 6 steps, only ControlNet is utilized to regulate the structural layout. Introducing IP-adapter in the 7th step guides appearance features. The final results reveal that, even if the Gaussian filter does not initially suppress redundant features, it remains effective in later stages during the generation of detailed textures. The above findings highlight that the impact of filtering on the diffusion process is intuitive, controllable, and predictable.
Figure 4: (a) shows the illustration of filter’s impact on the corresponding sampling inference stages in the static generative model and the dynamic generative model. (b) gives a comparison of the results obtained by applying filter to some works of generative models. The gray background represents traditional work based on GAN and AE architectures [16, 61, 55]. The yellow background represents work based on Diffusion [59, 56]. Comparing the results, we can intuitively see that filter operation has a more significant impact on diffusion models.
@@
In the early days, the fast style transfer algorithm [15, 28, 16] is based on CNN, AE, and GAN [10, 61, 55], and their fitting effect on the target data distribution depends on the design of the training process. During the training process, the model attempts to capture the statistical characteristics of the entire data distribution and obtain the maximum likelihood representation of the data distribution. When training is completed, the generated samples exist statically in the latent space. Therefore, we consider this type of model to be the static generative model.
@@
We noticed that prompts in pixel space may exhibit certain limitations in previous static generation models, but it can be well combined with dynamic generation models. As shown in Figure 4, we applied the same level of blur interference to the input data of various generative models. The results indicate that the impact of blur prompt on static generative models is relatively weak, while the outcomes of diffusion models show significant changes. We believe that for static generation models, using prompt on input data during the sampling inference stage will only have a slight impact on the mapping position of this input in the static distribution. And this position is close to the mapping result of the original image $x$. Therefore, prompts in pixel space will not significantly affect the generation results of static generation models. However, for diffusion models, as indicated by Equation 1, there is a predefined long-term dependency between the current time node image $x_{t}$ and the input image $x_{0}$, in the forward process. Therefore, the impact of using prompts on the input structure image $x_{0}$ will be executed again at each sampling, and the minor changes caused by prompts will participate in the entire generation process, thereby exerting a more significant influence on the generation results.
4 Our Approach
Figure 5: Our framework. The experiment uses ControlNet and IP-Adapter as the baseline and adds combined filtering operations as the expansion. We mapped low-level features in appearance images to global embeddings as $Cs$, concatenating them with SDM default text prompt embeddings $Ct$. The denoising generation processes these parts separately. A segment is managed by ControlNet, projecting latent distributions into a fused distribution controlled by high-level features that is $Cc$. The other part uses IP-Adapter for decoding and guiding low-level feature generation. Intermediate hidden state $x_{t-1}$ from both processes are weighted and summed every sampling time.
@@ 4 Our Approach
We propose the adoption of pixel-space methodology, FilterPrompt, to directly manipulate the frequency or distribution characteristics of specific image attributes, thereby influencing the subsequent expression levels of the representation. Our approach offers an intuitive and straightforward approach, and significantly saves computational overhead.